OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 3y – x = 12 x – 3y = 6

OpenStudy (anonymous):

Can you please show me the graphs?

OpenStudy (anonymous):

Or is the question just asking for the point of intersection between the two equations?

OpenStudy (danjs):

if you put them into standard form first, you can see something about those 2 equations

OpenStudy (danjs):

-x + 3y = 12 x - 3y = 6 notice

OpenStudy (anonymous):

3y – x = 12 x – 3y = 6 3y = x + 12 y = 1/3 x + 4

OpenStudy (danjs):

the coefficients on the x and y terms are a constant multiple of one to the othe requation multiply the first equation by -1

OpenStudy (danjs):

if this is the case, the lines are parallel

OpenStudy (anonymous):

x – 3y = 6 -3y = -x + 6 y = 1/3 x -2 and yup you are correct

OpenStudy (danjs):

Multiply first equation by -1 x - 3y = -12 x - 3y = 6

OpenStudy (danjs):

if the left side also becomes the same, then you just have one line kin two forms, just multiples of the equation

OpenStudy (danjs):

right side , sorry,

OpenStudy (anonymous):

That's right, I can see why you multiplied by -1

OpenStudy (anonymous):

Yea they're parallel same slope

OpenStudy (danjs):

i think that is what they are looking for from the looks of it, parallel , perpendicular, one solution intersection parallel lines = no solutions same line = infinit solutins intersecting, one soln

OpenStudy (danjs):

yes, for systems of linear equations, you can -change order of equations -multiply by constants -add equations together

OpenStudy (anonymous):

Yup

OpenStudy (danjs):

you can see , if you initially add the two equations together, you get 0 = 18 false, no solutions,no intersections, parallel

OpenStudy (anonymous):

thxs

OpenStudy (danjs):

welcome, anytime